40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16721     Entry time: Thu Mar 10 09:39:59 2022
Author: Jordan 
Type: Update 
Category: VAC 
Subject: Leak Testing of New Manual Gate Valve - Attempt #2 

This morning Chub and I leak checked the manual gate valve with the RGA and helium. There was no change in the helium partial pressure while spraying helium around the flanges, all looks good.

 

I also took a 100 AMU analog scan, after the filament had warmed up overnight and the plot was quite noisy even with the lowest scan speed. I recommend this unit go back to SRS for a filament replacement/recalibration. I am worried yesterday's "vent" of the RGA volume may have burned the filament. See the comparison of yesterday's analog scan to today's below.

Attachment 1: 3-09-2022.PNG  138 kB  | Hide | Hide all
3-09-2022.PNG
Attachment 2: 3-10-2022.PNG  122 kB  | Hide | Hide all
3-10-2022.PNG
ELOG V3.1.3-